MICROWAVE SPECTRA OF ^{18}O-TROPOLONE AND D-TROPOLONE

\underline{WEI LIN}, WALLACE C. PRINGLE, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459; THOMAS A. BLAKE, Pacific Northwest National Laboratory, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC05-76RLO 1830); JOHN C. KESKE, and DAVID F. PLUSQUELLIC, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, Maryland 20899-8441.

Tropolone is a seven-membered ring with an intramolecular hydrogen bond between the hydroxyl proton on C_{1} and the carbonyl oxygen on C_{2}. The proton tunnels between the two neighboring oxygen atoms. Microwave spectra of singly-substituted ^{18}O-tropolone and of singly-deuterated tropolone were measured to determine the effect of these two substitutions on the proton tunneling dynamics. The 0^{+} state of ^{18}O-tropolone and both the 0^{+} and 0^{−} states of Dtropolone were observed and assigned. Analysis of the spectra reveals complete quenching of the tunneling splitting of 18O-tropolone. The energy separation between the 0^{+} and 0^{−} states in D-tropolone was determined and is found to be significantly smaller than that of normal tropolone.